The following is a 3 x 3 two-way table:

According to this table: 

 a) A/P is a probability. b) N/P is a probability. c) F/H is a probability.
Answer #1

a) \frac{A}{P} is a JOINT probability

b)\frac{N}{P} is a MARGINAL probability

c)\frac{F}{H} is a CONDITIONAL probability
Answer #2

ornt Yn